OpenStudy (anonymous):

What's the simplified form of y^2+7y+12/y^2-2y-15

OpenStudy (anonymous):

Factorize both the quadratic equations first and then divide them. Can you do it ?

OpenStudy (anonymous):

Do what?

OpenStudy (anonymous):

Factorize the equations.

OpenStudy (jhannybean):

\[\large \frac{y^{2}+7y+12}{y^{2}-2y-15}\]factor out both the numerator and denominator. numerator : what two numbers add to give 7 and multiply to give 12? denominator: what to numbers add to give -2 and multiply to give -15?

OpenStudy (anonymous):

numerator: 3 and 4 denominator: 3 and -5

OpenStudy (jhannybean):

correct :) \[\large \frac{(x+3)(x+4)}{(x-5)(x+3)}\] Now is there any like terms we can cancel out?

OpenStudy (anonymous):

x+3

OpenStudy (jhannybean):

good! \[\large \frac{\cancel{(x+3)}(x+4)}{(x-5)\cancel{(x+3)}}\]

OpenStudy (jhannybean):

What does that leave us with?

OpenStudy (anonymous):

x+4/x-5

OpenStudy (jhannybean):

You got it :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Using the Discriminant method, you'll get numerator=(y+3)(y+4) And denominator=(y-5)(y+3) So the answer will be (y+4)/(y-5)

OpenStudy (jhannybean):

\[\large \frac{y^{2}+7y+12}{y^{2}-2y-15} =\large \frac{(x+3)(x+4)}{(x-5)(x+3)}=\large \frac{\cancel{(x+3)}(x+4)}{(x-5)\cancel{(x+3)}}= \frac{x+4}{x+3}\]

OpenStudy (anonymous):

A little bit of correction to the above comment. Ans= (y+4)/(y-5) :) Y+3 just got cancelled out.

OpenStudy (jhannybean):

Somehow my y's magically changed to x's. lol

OpenStudy (anonymous):

I wasn't pointing towards X or Y. You wrote an incorrect denominator in the answer. That's what I was talking about.

OpenStudy (jhannybean):

Oh that lol.-_- I see it